OpenStudy (luigi0210):

Another Limit problem:

OpenStudy (luigi0210):

\[\lim_{x \rightarrow 0} \frac{x}{\frac{1}{6}+\frac{1}{x-6}}\]

OpenStudy (yrelhan4):

You just need to simplify it. Your main objective is eliminating 'x'. Simplifying the expression will do that for you.

OpenStudy (jdoe0001):

\(\bf \lim_{x \rightarrow 0} \cfrac{x}{\frac{1}{6}+\frac{1}{x-6}}\\ \cfrac{ x }{ \frac{x-6+6}{6(x-6)} } \implies \cfrac{ x }{ \frac{x}{6(x-6)} } \implies \cfrac{\cancel{x}}{1} \times \cfrac{6(x-6)}{\cancel{x}} \)

OpenStudy (anonymous):

\[\frac{x}{\frac{x}{6(x-6)}}=6(x-6)\]

OpenStudy (dls):

@yrelhan4 hogaya rehhne de. easy tha :P

OpenStudy (yrelhan4):

I thought you were just supposed to give hints and let the asker work it out.

OpenStudy (luigi0210):

Apprently not @yrelhan4 xD Anyways thanks guys :)

OpenStudy (dls):

ain't anyone got chance to showoff their skills

OpenStudy (luigi0210):

* apparently

OpenStudy (jdoe0001):

hehe, well, one can say it was just a bit longer hint heheh anyhow, we didn't type the actual value :)

OpenStudy (dls):

lol

OpenStudy (luigi0210):

But anyways, i understood what they did.